The relationship between pressure and temperature, when volume and moles of a gas are held constant, is: P/T = k. We could say then, that:

If the temperature is doubled, the pressure must be halved.

If the temperature is halved, the pressure must also be halved.

If the temperature is halved, the pressure must double.

If the temperature is doubled, the pressure must remain the same.
1
Expert's answer
2017-05-23T05:55:30-0400
Answer:
If the temperature is halved, the pressure must also be halved.

Solution

According to the formula given, the proportion between pressure and temperature remains constant. This means that when one parameter doubles, the other should also double:
p/T=(2×p)/(2×T)=(0.5×p)/(0.5×T)=k